Complete the ordered pairs for the function h(a)=4a-5
a. (2,?) b. (11,?)

Answer by funmath(2933)

Complete the ordered pairs for the function h(a)=4a-5
a. (2,?)
Let a=2 and solve for h(2):
h(2)=4(2)-5
h(2)=8-5
h(2)=3
The ordered pair is: (2,3)
b. (11,?)
Let a=11 and solve for h(11):
h(11)=4(11)-5
h(11)=44-5
h(11)=39
The ordered pair is: (11,39)
Happy Calculating!!!!